# LTGC: Long-tail Recognition via Leveraging LLMs-driven Generated Content

### 09 Mar 2024 v1

@@ 1 Introduction
In the real world, data often exhibits a long-tailed distribution, posing significant challenges for computer vision recognition [56, 46]. These challenges include (1) Class Imbalance: within the dataset, some classes (termed ”head” classes) are abundantly represented, while others (termed ”tail” classes) have few samples. This imbalanced distribution during training may cause the model to focus more on the head classes, neglecting the tail classes [44]. (2) Tail Data Scarcity: Data scarcity refers to tail classes having an extremely limited number of samples, which lack diversity and are insufficient to effectively train a model [24, 8]. It prevents a model’s ability to learn the feature invariant, which is necessary to recognize these categories correctly [35].
@@
To address these challenges, numerous approaches have emerged within the field of long-tailed recognition, such as re-sampling [5, 19], loss re-weighting [23, 26, 12, 6, 50, 1, 29, 27, 41], ensemble learning [49, 44, 57], decoupling [20], and contrastive learning [42, 61, 11, 25]. Their primary objective is to balance the decision boundaries and parameter weights of the model to improve long-tail recognition. However, these methods encounter bottlenecks due to the scarcity and limited diversity features of tail-class data [24]. In order to obtain more diverse representations for tail classes, two types of methods appear: Some methods increase tail-class diversity through data augmentation [9, 24]. Others increase tail-class diversity by transferring features from related classes [8] or large pre-trained models (e.g., CLIP [32]). These transfer learning methods [14, 28, 36] based on CLIP have recently shown great potential in boosting long-tail recognition. The LPT [14] fine-tunes pre-trained models to adapt to target datasets. The VL-LTR [36] leverages text features from CLIP to augment the learning of image features. However, these methods find it hard to obtain the correct and desired knowledge for tail categories. For example, in the real-world data iNauralist 2018 [17], the category Aquilegia Pubescens is characterized only by white or pale yellow colors and spurred shapes. If we transfer similar category features or blindly augment semantic information, such as red color semantics, this category will become more confusing.
@@
Recently, Large Language Models (LLMs) (e.g., ChatGPT) and Large Multimodal Models (LMMs) (e.g., GPT-4V (ChatGPT with Vision [52]) and Minigpt4 [60, 7]) due to their wealth of implicit knowledge, have been leveraged for a variety of downstream tasks, such as robot task plans [34], open-set object recognition [30], automated robot learning [45] and various visual reasoning tasks [52, 53]. However, due to the bias of imbalanced training data, LMMs perform poorly on some complex tasks [10]. We also find this dilemma in long-tail recognition as shown in Fig. 1. When we asked LMMs about an image in the category of Trogon rufus: ”What species is in the image? ”, we did not get the expected answer. In contrast, when we ask some easy questions about this image: ”Please describe the image.” or ”Please describe the distinctive features of Trogon Rufus”, we get the desired answer. This suggests that although large models underperform in long-tail recognition (e.g., it is difficult to align complex image features with labels.), they still contain desired implicit knowledge (e.g., correctly describing images and providing the knowledge of species). Inspired by this, we aim to leverage the rich implicit knowledge of large models to address the challenges of long-tailed data scarcity and perform long-tailed recognition well. Nevertheless, effectively leveraging large models to encounter these challenges is not trivial: (1) The implicit nature of knowledge within large models makes it difficult to extract the desired knowledge that facilitates long-tail recognition tasks. (2) Due to the domain gap between the generated data and the original data [38], effectively using these hybrid data remains an open issue.
@@
To this end, we propose a novel Long-Tail recognition framework via Generated Content, denoted as LTGC, which is illustrated in Fig. 2. Our LTGC aims to leverage the capabilities of large models for generating explicitly diverse content tailored to the long-tail classes and incorporate novel designs to enhance long-tail recognition. Firstly, motivated by the wealth of implicit knowledge in large models and the fact that texts are more controllable, we employ them to expand the tail classes. Specifically, to produce more diverse and controllable tail data, we follow the rule of analyzing existing features before generating absent features. To achieve this, LTGC utilizes LMMs to analyze images already in the tail classes to obtain the existing tail-class descriptions list. Then LTGC inputs the existing descriptions list to LLMs to obtain the desired extended tail-class descriptions list for images that are absent from the existing tail classes. Secondly, LTGC utilizes the text-to-image (T2I) model (e.g., DALL-E [33]) to transform these textual descriptions into images. Moreover, inspired by the benefits of chain-of-thought [48], we propose a self-reflection and iterative evaluation module for this process to ensure the diversity and quality of the generated content. Finally, inspired by the advantages of Mixup [54] in merging different image domains, we propose the BalanceMix module to address domain shifts of generated images for the fine-tuning process.
@@ 2 Related Work
Large Models. Recently, several large models have emerged, including large language models (e.g., ChatGPT [52]), large multi-modal models (e.g., GPT-4 (ision) [52], MiniGPT [60]) and large generated models (e.g., DALL-E [33]). These expansive models are repositories of extensive knowledge and have undergone scrutiny in diverse applications, such as robot task plans [34], open-set object recognition [30], contextual object detection [53], and video generation [22]. In this work, we leverage the rich implicit knowledge of these large models to handle the long-tail recognition challenge.
@@ 3 Method
To perform long-tail recognition well, the challenge is making the model learn diverse representations from tail categories and address imbalances across all categories. In this paper, we propose a novel long-tail recognition framework, LTGC, to handle these challenges. As depicted in Fig. 2, LTGC extracts implicit knowledge from various off-the-shelf large models to generate and cyclically assess the quality and diversity of tail classes (Sec. 3.1). Furthermore, LTGC proposes the BalanceMix module to facilitating the fine-tuning process with the generated and original datasets (Sec. 3.2). These modules are described in detail below.
@@ 3.1.2 Obtaining Extended Tail-class Descriptions List
In addition, in order for these generated descriptions to better complement each tail class, we encourage LLMs to generate descriptions of sufficient number and diversity. To achieve this, we introduce a self-reflection module in this process, aiming to guide LMMs in rethinking if there are any features or scenes that are missed or repeated. It includes two key designs: a number-checking module and a repetition-checking module. (1) The number-checking module aims to guide LMMs in rethinking if there are other features or scenes that are missed. To achieve this, after posing the initial [Prompt 2] for class $y$, we update the extended descriptions list and re-ask LLMs the [Prompt 2] question, incorporating the newly acquired list. For each class $y$, this iterative process of number-checking continues until a maximum number $K_{y}$ of the tail class is achieved, where $K_{y}$ = $M_{y}$+$N_{y}$ and $N_{y}$ is the number of generated descriptions for class $y$. (2) The repetition-checking module aims to guide LMMs in rethinking if there are other features or scenes that are repeated at the end of the number-checking iteration. Specifically, we input the extended descriptions list and the following [Prompt 3] of each class $y$ for LLMs’ repetition checking: ”Please exclude any repetitive {distinctive features} and {specific scenes} for class $[y]$ in this descriptions list.” After that, LLM will filter out the ones with repeated features based on this prompt and return a new list of descriptions. Through the implementation of this two-step process and LLMs’ self-reflection module, we have obtained a reduction in repeated descriptions and an increase in the diversity of the tail-class descriptions for each class.
@@ 3.1.3 Transform Descriptions to Images
where $n\in N$ denotes the $n$-th generated descriptions for class $y$, and $i_{n}^{y}$ denotes the image generated by T2I model based on $d_{n}^{y}$. However, some generated images may not hold a high enough quality to accurately represent their desired classes, as the lower-quality images. As shown in Fig. 7, lower-quality images will output poorly distinctive features, resulting in more confusion for the class. Therefore, using these generated images directly in training may result in disrupting the model’s prediction of the tail classes.
@@
where $\text{Encoder}_{\text{vis}}$ denotes the CLIP’s visual encoder, and $\text{Encoder}_{\text{text}}$ denotes the CLIP’s text encoder. The image $i_{n}^{y}$ is detected as lower quality and filtered out if its similarity score $\mathcal{S}$ is below a threshold $\mu$. 2) Refinement. Next, also inspired by chain-of-thought [48], we propose the following prompt for refine its descriptions. Specifically, if $i_{n}^{y}$ is identified as a lower quality image, we prompt LLMs to refine its corresponding description $d_{n}^{y}$ to more accurately represent the intended class $y$, drawing on feedback provided by CLIP, i.e., [Prompt 5]: ”This description $d_{n}^{y}$ doesn’t seem to be representative of the class $[y]$. Could you refine it to enhance the distinctive features of class $[y]$?” 3) Re-generation. Finally, the image $i_{n}^{y}$ is regenerated through T2I model based on the refined textual description.
@@ 4 Experiments
Method Implementation. In our LTGC, we incorporate diverse and specialized knowledge from the off-the-shelf large models. Specifically, for LMM, we use the GPT-4V (ision) [52] version of ChatGPT. For LLM, we use the GPT-4 version of ChatGPT. For T2I, we use DAll-E [33]. For the pre-trained CLIP [32], we use ViT-B/32 [15] for its visual encoder and the transformer architecture described in [31] for its text encoder. In LLM’s self-reflection module, we set the maximum number $K_{y}$ to 100, 300, and 800 for iNaturalist 2018, ImageNet-LT, and Place-LT, respectively. In iterative evaluation module, the threshold $\mu$ is set at 0.8 for ImageNet-LT and Place-LT, and at 0.6 for iNaturalist. In Appendix, we provide a detailed ablation and discussion on the choice of LMM and LLM models, as well as parameters used in the self-reflection process.
@@ 4.2 Comparisons with SOTA on Benchmarks
Results on Places-LT. Tab. 1 shows that compared to other CLIP variant methods, LTGC achieves 54.1% and 52.1% in terms of overall accuracy and few-shot accuracy, respectively surpassing the LPT [14] by 4.0% and 5.2%. Even compared with VL-LTR [36] and RAC [28], which are CLIP-based long-tail methods with extra data in training and testing, our LTGC achieves the same overall accuracy while obtaining higher few-shot accuracy.
Results on iNaturalist 2018. Finally, we explore LTGC on a large-scale and fine-grained dataset, iNaturalist 2018. Tab. 2 presents the quantitative results. LTGC leverages the rich knowledge of LLMs and significantly outperforms traditional deep learning approaches for long-tail recognition. In addition, LTGC attains an overall accuracy of 82.5% and a few-shot accuracy of 82.6%, outperforming all existing SOTA methods based on CLIP. Especially, LTGC also surpasses the retrieval augmented method, RAC [28], by 2.3%. These results demonstrate that LTGC can also handle large-scale, long-tailed data with only prompt tuning and achieve comparable accuracy.
@@
In the experiments, our LTGC employs GPT-4V (ision) to describe a given image to obtain a text-based feature description. A straightforward way for long-tail recognition is by querying LMMs for the category of the given image. To evaluate the ability of these methods to recognize the long tail images, we constructed the following baselines of LMMs for comparison: MiniGPT4 [60], MiniGPT4-v2 [7], LENS [3], and GPT-4V (ision) [52]. To perform a fair evaluation, we provide the label list [Class 1, Class 2, …, Class Y] for each dataset, where $Y$ is the number of classes. The results presented in Tab. 3, show that our method significantly outperforms the baselines that directly query the LMMs. More implementation details are discussed in Appendix.
@@ 4.4 Analysis and Ablation Study
Effectiveness of the iterative evaluation module. To guarantee the accurate representation of the desired classes by the images produced via T2I, we have integrated an iterative evaluation module within our architecture for progressive refinement of images. To assess the effectiveness of this module, we contrasted it with three distinct image generation strategies: 1) w/o iterative evaluation: the images are fed directly into our framework’s training process without any preliminary detection or refinement. 2) Detection and exclusion: the CLIP model evaluates the generated images, selectively forwarding only the ones that align closely with the intended class criteria to the training phase. Images that fail to meet the detection threshold are excluded, bypassing the refinement step entirely. As shown in Tab. 4, the performance of the two variants is worse than our method. This suggests that our proposed iterative evaluation module incorporating filtering and refinement of the design is more effective.
@@
Figure 7: The visualization of the images generated before and after passing the iterative evaluation module. The top row displays images are filtered out, while the bottom row shows images regenerated by T2I after refining their corresponding descriptions. More visualizations are in Appendix.
Effectiveness of the BalanceMix module. To evaluate the effectiveness of the two proposed designs, we conducte experiments on three different variants: 1) w/o BalanceMix: This variant is directly fine-tuned using the generated data and the original training data. 2) Balanced Sample: This variant utilizes generated data to perform balanced sampling [47] for training. 3) Mixup: This variant employs generated data with Mixup [54] without performing balanced sampling. The results are shown in Tab. 5, demonstrating that our method significantly outperforms the other variants. Although our self-reflection and iterative evaluation modules already ensure the diversity and high quality of the generated images, there still exists a domain gap between the generated and original data [38]. This domain gap could exacerbate the long-tail effect on the test set (discussed in Appendix). A simple balanced sample [47] approach fails to solve the class imbalance caused by the mixing of original and generated data, while the Mixup [54] alone cannot address inter-class imbalances. To this end, our BalanceMix module combines the strengths of both methods, making the generated data well-suited for our framework.

### 13 Mar 2024 v3

@@ 1 Introduction
In the real world, data often exhibits a long-tailed distribution, posing significant challenges for computer vision recognition [zhang2023deep, 46]. These challenges include (1) Class Imbalance: within the dataset, some classes (termed ”head” classes) are abundantly represented, while others (termed ”tail” classes) have few samples. This imbalanced distribution during training may cause the model to focus more on the head classes, neglecting the tail classes [44]. (2) Tail Data Scarcity: Data scarcity refers to tail classes having an extremely limited number of samples, which lack diversity and are insufficient to effectively train a model [24, 8]. It prevents a model’s ability to learn the feature invariant, which is necessary to recognize these categories correctly [35].
3 Refer to caption
@@
To address these challenges, numerous approaches have emerged within the field of long-tailed recognition, such as re-sampling [5, 19], loss re-weighting [23, 26, 12, 6, 50, 1, 29, 27, 41], ensemble learning [49, 44, 57], decoupling [20], and contrastive learning [42, 61, 11, 25]. Their primary objective is to balance the decision boundaries and parameter weights of the model to improve long-tail recognition. However, these methods encounter bottlenecks due to the scarcity and limited diversity features of tail-class data [24]. To obtain more diverse representations for tail classes, two types of methods appear: Some methods increase tail-class diversity through data augmentation [9, 24, mixup, zhao2022mixpro]. Others increase tail-class diversity by transferring features from related classes [8] or large pre-trained models (e.g., CLIP [32]). These transfer learning methods [14, 28, 36] based on CLIP have recently shown great potential in boosting long-tail recognition. The LPT [14] fine-tunes pre-trained models to adapt to target datasets. The VL-LTR [36] leverages text features from CLIP to augment the learning of image features. However, these methods find it hard to obtain the correct and desired knowledge for tail categories. For example, in the real-world data iNauralist 2018 [17], the category Aquilegia Pubescens is characterized only by white or pale yellow colors and spurred shapes. If we transfer similar category features or blindly augment semantic information, such as red color semantics, this category will become more confusing.
@@
Recently, Large Language Models (LLMs) (e.g., ChatGPT) and Large Multimodal Models (LMMs) (e.g., GPT-4V (ChatGPT with Vision [52]) and Minigpt4 [60, 7]) due to their wealth of implicit knowledge, have been leveraged for a variety of downstream tasks, such as robot task plans [34], open-set object recognition [30], automated robot learning [45] and various visual reasoning tasks [52, 53, zhu2023llafs]. However, due to the bias of imbalanced training data, LMMs perform poorly on some complex tasks [10]. We also find this dilemma in long-tail recognition as shown in Fig. 1. When we asked LMMs about an image in the category of Trogon rufus: ”What species is in the image? ”, we did not get the expected answer. In contrast, when we ask some easy questions about this image: ”Please describe the image.” or ”Please describe the distinctive features of Trogon Rufus”, we get the desired answer. This suggests that although large models underperform in long-tail recognition (e.g., it is difficult to align complex image features with labels.), they still contain desired implicit knowledge (e.g., correctly describing images and providing the knowledge of species). Inspired by this, we aim to leverage the rich implicit knowledge of several large models to address the challenges of long-tailed data scarcity and perform long-tailed recognition well. Nevertheless, effectively leveraging large models to encounter these challenges is not trivial: (1) The implicit nature of knowledge within large models makes it difficult to extract the desired knowledge that facilitates long-tail recognition tasks. (2) Due to the domain gap between the generated data and the original data [38], effectively using these hybrid data remains an open issue.
@@
To this end, we propose a novel Long-Tail recognition framework via Generated Content, denoted as LTGC, which is illustrated in Fig. 2. Our LTGC aims to leverage the capabilities of large models for generating explicitly diverse content [foo2023aigenerated] tailored to the long-tail classes and incorporate novel designs to enhance long-tail recognition. Firstly, motivated by the wealth of implicit knowledge in large models and the fact that texts are more controllable, we employ them to expand the tail classes. Specifically, to produce more diverse and controllable tail data, we follow the rule of analyzing existing features before generating absent features. To achieve this, LTGC utilizes LMMs to analyze images already in the tail classes to obtain the existing tail-class descriptions list. Then LTGC inputs the existing descriptions list to LLMs to obtain the desired extended tail-class descriptions list for images that are absent from the existing tail classes. Secondly, LTGC utilizes the text-to-image (T2I) model (e.g., DALL-E [33]) to transform these textual descriptions into images. Moreover, inspired by the benefits of chain-of-thought [48], we propose a self-reflection and iterative evaluation module for this process to ensure the diversity and quality of the generated content. Finally, inspired by the advantages of Mixup [54] in merging different image domains, we propose the BalanceMix module to address domain shifts of generated images for the fine-tuning process.
@@ 2 Related Work
Large Models. Recently, several large models have emerged, including large language models (e.g., ChatGPT [52]), large multi-modal models (e.g., GPT-4 (ision) [52], MiniGPT [60]) and large generated models (e.g., DALL-E [33]). These expansive models are repositories of extensive knowledge [foo2023aigenerated] and have undergone scrutiny in diverse applications, such as robot task plans [34], open-set object recognition [30], contextual object detection [53], video generation [22], and few-shot segmentation [zhu2023llafs]. In this work, we leverage the rich implicit knowledge of these large models to handle the long-tail recognition challenge.
@@ 3 Method
To perform long-tail recognition well, the challenge is making the model learn diverse representations from tail categories and address imbalances across all categories. In this paper, we propose a novel long-tail recognition framework, LTGC, to handle these challenges. As depicted in Fig. 2, LTGC employ implicit knowledge from various off-the-shelf large-scale models to generate and iteratively assess the quality and diversity of tail classes (Sec. 3.1). Furthermore, LTGC proposes the BalanceMix module to facilitating the fine-tuning process with the generated and original datasets (Sec. 3.2). These modules are described in detail below.
@@ 3.1.2 Obtaining Extended Tail-class Descriptions List
In addition, in order for these generated descriptions to better complement each tail class, we encourage LLMs to generate descriptions of sufficient number and diversity. To achieve this, we introduce a self-reflection module in this process, aiming to guide LMMs in rethinking if there are any features or scenes that are missed or repeated. It includes two key designs: a number-checking module and a repetition-checking module. (1) The number-checking module aims to guide LMMs in rethinking whether there are other missing features or scenarios. To achieve this, after posing the initial [Prompt 2] for class $y$, we update the extended descriptions list and re-ask LLMs the [Prompt 2] question, incorporating the newly acquired list. For each class $y$, this iterative process of number-checking continues until a maximum number $K_{y}$ of the tail class is achieved, where $K_{y}$ = $M_{y}$+$N_{y}$ and $N_{y}$ is the number of generated descriptions for class $y$. (2) The repetition-checking module aims to guide LMMs in rethinking if there are other features or scenes that are repeated at the end of the number-checking iteration. Specifically, we input the extended descriptions list and the following [Prompt 3] of each class $y$ for LLMs’ repetition checking: ”Please exclude any repetitive {distinctive features} and {specific scenes} for class $[y]$ in this descriptions list.” After that, LLM will filter out the ones with repeated features based on this prompt and return a new list of descriptions. Through the implementation of this two-step process and LLMs’ self-reflection module, we have obtained a reduction in repeated descriptions and an increase in the diversity of the tail-class descriptions for each class.
@@ 3.1.3 Transform Descriptions to Images
where $n\in N$ denotes the $n$-th generated descriptions for class $y$, and $i_{n}^{y}$ denotes the image generated by T2I model based on $d_{n}^{y}$. However, some generated images may not be of sufficiently high quality to accurately represent their desired classes, as the lower-quality images. As shown in Fig. 7, lower-quality images will output poorly distinctive features, resulting in more confusion for the class. Therefore, using these generated images directly in training may result in disrupting the model’s prediction of the tail classes.
@@
where $\text{Encoder}_{\text{vis}}$ denotes the CLIP’s visual encoder, and $\text{Encoder}_{\text{text}}$ denotes the text encoder of the CLIP . The image $i_{n}^{y}$ is detected as lower quality and filtered out if its similarity score $\mathcal{S}$ is below a threshold $\mu$. 2) Refinement. Next, also inspired by chain-of-thought [48], we propose the following prompt for refine its descriptions. Specifically, if $i_{n}^{y}$ is identified as a lower quality image, we prompt LLMs to refine its corresponding description $d_{n}^{y}$ to more accurately represent the intended class $y$, drawing on feedback provided by CLIP, i.e., [Prompt 5]:
”This description $d_{n}^{y}$ doesn’t seem to be representative of the class $[y]$. Could you refine it to enhance the distinctive features of class $[y]$?” 3) Re-generation. Finally, the image $i_{n}^{y}$ is regenerated by the T2I model according to the improved textual description.
@@ 4 Experiments
Method Implementation. In our LTGC, we incorporate diverse and specialized knowledge from the off-the-shelf large models. Specifically, for LMM, we use the GPT-4V (ision) [52] version of ChatGPT. For LLM, we use the GPT-4 version of ChatGPT. For T2I, we use DAll-E [33]. For the pre-trained CLIP [32], we use ViT-B/32 [15] for its visual encoder and the transformer architecture described in [31] for its text encoder. In LLM’s self-reflection module, we set the maximum number $K_{y}$ to 100, 300, and 800 for iNaturalist 2018, ImageNet-LT, and Place-LT, respectively. In the iterative evaluation module, the threshold $\mu$ is set at 0.8 for ImageNet-LT and Place-LT, and at 0.6 for iNaturalist. In Appendix, we provide a detailed ablation and discussion on the choice of LMM and LLM models, as well as parameters used in the self-reflection process.
@@ 4.2 Comparisons with SOTA on Benchmarks
Results on Places-LT. Tab. 1 shows that compared to other CLIP variant methods, LTGC achieves 54.1% and 52.1% in terms of overall accuracy and few-shot accuracy, respectively surpassing the LPT [14] by 4.0% and 5.2%. Even compared with VL-LTR [36] and RAC [28], which have extra data in training and testing, our LTGC achieves remarkable results.
Results on iNaturalist 2018. Finally, we explore LTGC on a large-scale and fine-grained dataset, iNaturalist 2018. Tab. 2 presents the quantitative results. LTGC leverages the rich knowledge of LLMs and significantly outperforms traditional deep-learning approaches for long-tail recognition. In addition, LTGC attains an overall accuracy of 82.5% and a few-shot accuracy of 82.6%, outperforming all existing SOTA methods based on CLIP. In particular, LTGC also surpasses the retrieval augmented method, RAC [28], by 2.3%.
@@
In the experiments, our LTGC employs GPT-4V (ision) to describe a given image to obtain a text-based feature description. A straightforward way for long-tail recognition is by querying LMMs for the category of the given image. To evaluate the ability of these methods to recognize the long tail images, we constructed the following baselines of LMMs for comparison: MiniGPT4 [60], MiniGPT4-v2 [7], LENS [3], and GPT-4V (ision) [52]. To perform a fair evaluation, we provide the label list [Class 1, Class 2, …, Class Y] for each dataset, where $Y$ is the number of classes.
The results presented in Tab. 3, show that our method significantly outperforms the baselines that directly query the LMMs. More implementation details are discussed in Appendix.
@@ 4.4 Analysis and Ablation Study
Effectiveness of the iterative evaluation module. To guarantee the accurate representation of the desired classes by the images produced via T2I, we have integrated an iterative evaluation module within our architecture for the progressive refinement of images. To assess the effectiveness of this module, we contrasted it with three distinct image generation strategies: 1) w/o iterative evaluation: the images are fed directly into our framework’s training process without any preliminary detection or refinement. 2) Detection and exclusion: the CLIP model evaluates the generated images, selectively forwarding only the ones that align closely with the intended class criteria to the training phase. Images that fail to meet the detection threshold are excluded, bypassing the refinement step entirely. As illustrate in Tab. 4, the performance of the two variants is worse than our method. This suggests that our proposed iterative evaluation module incorporating filtering and refinement of the design is more effective.
@@
Figure 7: The visualization of the images generated before and after passing the iterative evaluation module. The top row displays images that were filtered out, while the bottom row shows images regenerated by T2I after refining their corresponding descriptions. More visualizations are in Appendix.
Effectiveness of the BalanceMix module. To evaluate the effectiveness of the two proposed designs, we conduct experiments on three different variants: 1) w/o BalanceMix: This variant is directly fine-tuned using the generated data and the original training data. 2) Balanced Sample: This variant utilizes generated data to perform balanced sampling [47] for training. 3) Mixup: This variant employs generated data with Mixup [54] without performing balanced sampling. The results are shown in Tab. 5, demonstrating that our method significantly outperforms the other variants. Although our self-reflection and iterative evaluation modules already ensure the diversity and high quality of the generated images, there still exists a domain gap between the generated and original data [38]. This domain gap could exacerbate the long-tail effect on the test set (discussed in Appendix). A simple balanced sample [47] approach fails to solve the class imbalance caused by the mixing of original and generated data, while the Mixup [54] alone cannot address inter-class imbalances. To this end, our BalanceMix module combines the strengths of both methods, making the generated data well-suited for our framework.